# A Flow-Based Model for Conditional and Probabilistic Electricity Consumption Profile Generation and Prediction

### v2

@@ Abstract
Residential Load Profile (RLP) generation and prediction are critical for the operation and planning of distribution networks, particularly as diverse low-carbon technologies are increasingly integrated. This paper introduces a novel flow-based generative model, termed Full Convolutional Profile Flow (FCPFlow), which is uniquely designed for both conditional and unconditional RLP generation, and for probabilistic load forecasting. By introducing two new layers–the invertible linear layer and the invertible normalization layer–the proposed FCPFlow architecture shows three main advantages compared to traditional statistical and contemporary deep generative models: 1) it is well-suited for RLP generation under continuous conditions, such as varying weather and annual electricity consumption, 2) it shows superior scalability in different datasets compared to traditional statistical models, and 3) it also demonstrates better modeling capabilities in capturing the complex correlation of RLPs compared with deep generative models.
@@ II Modeling of Residential Load Profiles
where $x_{i,t}$ is the active power consumption of $t$-th time step, $\mathbf{x_{i}}=(x_{1,i},...,x_{T,i})$ represents one RLP. An unconditional deep generative model (e.g., GAN) can be trained to generate RLPs. Such a generative model can be expressed as $G_{\theta}(\mathbf{z_{i}})=\mathbf{x_{i}}$, where $G(\cdot)$ is the generative model which maps $\mathbf{z}$ to $\mathbf{x}$, $\theta$ is the learnable parameters, $\mathbf{z}\sim\pi(\mathbf{z})$, and $\pi(\mathbf{z})$ can be any simple distribution such as Gaussian distribution. Then, a conditional deep generative model can be expressed as
@@ III-A Flow Based Models
The basic structure of conditional flow-based models is shown in Fig. 1, where $f$ (usually constructed by neural networks (NNs)) is essentially the generator $G_{\theta}(\cdot)$ in (2). During training, function $f^{-1}$ is learned to transform input data $\mathbf{x}$ (with condition $\mathbf{c}$) into $\mathbf{z}$ which follows a standard Gaussian distribution. Since the function $f^{-1}$ is invertible, once $f^{-1}$ is trained, its inverse $f$ is used to take random samples $\mathbf{z}$ (with condition $\mathbf{c}$) and generate $\mathbf{x}$. In flow-based models, function $f$ (or $f^{-1}$) is usually constructed by stacking multiple invertible transformations $f_{i}$, meaning $f=f_{1}\circ f_{2}...\circ f_{K}$ and $f^{-1}=f^{-1}_{K}\circ f^{-1}_{K-1}...\circ f^{-1}_{1}$. Fig 2 demonstrates how this stacked function transforms a simple Gaussian distribution into data distribution $p_{K}(\mathbf{x|c})$ and vice versa. By stacking invertible transformations $f_{i}$, the modeling capability of $f$ is also increased, enabling function $f$ to simulate more complex data distribution.
@@
Figure 1: Structure of flow-based models, where $\mathbf{x}$ is the input data, $\mathbf{c}$ is the condition corresponding to data, $\mathbf{z}$ is a latent variable following Gaussian distribution, and $f$ is a bijective function $f:Z\rightarrow X$ which is usually constructed by NNs.
@@ III-B Combining Coupling Layer
$\mathbf{z_{1}}=exp(s_{1}(\mathbf{x_{1},c}))\odot\mathbf{x_{2}}+t_{1}(\mathbf{x_{1},c})$ (9)
$\mathbf{(z_{1},c)}=Combine(\mathbf{z_{1},c})$ (10)
$\mathbf{z_{2}}=exp(s_{2}(\mathbf{z_{1},c}))\odot\mathbf{x_{1}}+t_{2}(\mathbf{z_{1},c})$ (11)
$\mathbf{z}=Combine(\mathbf{z_{1},z_{2}}),$ (12)
@@
Figure 3: The conditional transformation architecture of combining coupling layer $f_{ccl}$ and its inverse. $t$ and $s$ are the neural networks that can be expressed as $t(\cdot)$ and $s(\cdot)$, both $s$ and $t$ reduce the input dimensions as demonstrated in the figure.
[…]
@@
$\mathbf{x_{1}}=(\mathbf{z_{2}}-t_{2}(\mathbf{z_{1},c}))/exp(s_{2}(\mathbf{z_{1},c}))$ (15)
$\mathbf{(x_{1},c)}=Combine(\mathbf{x_{1},c})$ (16)
$\mathbf{x_{2}}=(\mathbf{z_{1}}-t_{1}(\mathbf{x_{1},c}))/exp(s_{1}(\mathbf{x_{1},c}))$ (17)
$\mathbf{x}=Combine(\mathbf{x_{1},x_{2}}).$ (18)
@@ IV Proposed Model: Full Convolutional Profile Flow
Figure 4: FCPFlow architecture is structured to process data through a series of FCPFlow blocks. $\mathbf{x}$ and $\mathbf{\hat{x}}$ are the input RLPs and generated RLPs, respectively. $\mathbf{z}$ and $\mathbf{\hat{z}}$ are the generated and sampled latent variables, respectively. $\beta$ and $\gamma$ are parameters in $f_{norm_{i}}$ introduced in Sec. IV-A, $W$ is a matrix that linearly transforms the input vector into another vector, $W$ is introduced in Sec. IV-C.
@@ IV-C Maximum Likelihood Estimation of FCPFlow
NL 60 minutes 27,757
AUS[35] 30 minutes 10,000
UK 30 minutes 10,000
USA[36] 15 minutes 9,110
UK weather[37] 30 minutes 10,000
@@
NL 60 minutes 365
UK 30 minutes 365
USA[36] 15 minutes 365
@@ VII Simulation Results for RLP Probabilistic Prediction
NL dataset (60 minutes resolution)
@@
UK dataset (30 minutes resolution)
@@
USA dataset (15 minutes resolution)

### v3

@@ Abstract
Residential Load Profile (RLP) generation and prediction are critical for the operation and planning of distribution networks, especially as diverse low-carbon technologies (e.g., photovoltaic and electric vehicles) are increasingly adopted. This paper introduces a novel flow-based generative model, termed Full Convolutional Profile Flow (FCPFlow), which is uniquely designed for both conditional and unconditional RLP generation, and for probabilistic load forecasting. By introducing two new layers–the invertible linear layer and the invertible normalization layer–the proposed FCPFlow architecture shows three main advantages compared to traditional statistical and contemporary deep generative models: 1) it is well-suited for RLP generation under continuous conditions, such as varying weather and annual electricity consumption, 2) it demonstrates superior scalability in different datasets compared to traditional statistical models, and 3) it also demonstrates better modeling capabilities in capturing the complex correlation of RLPs compared with deep generative models.
@@ II Modeling of Residential Load Profiles
where $x_{t,i}$ is the active power consumption of $t$-th time step, $\mathbf{x_{i}}=(x_{1,i},...,x_{T,i})$ represents one RLP. An unconditional deep generative model (e.g., GAN) can be trained to generate RLPs. Such a generative model can be expressed as $G_{\theta}(\mathbf{z_{i}})=\mathbf{x_{i}}$, where $G(\cdot)$ is the generative model which maps $\mathbf{z}$ to $\mathbf{x}$, $\theta$ is the learnable parameters, $\mathbf{z}\sim\pi(\mathbf{z})$, and $\pi(\mathbf{z})$ can be any simple distribution such as Gaussian distribution. Then, a conditional deep generative model can be expressed as
@@ III-A Flow Based Models
The basic structure of conditional flow-based models is shown in Fig. 1, where $f$ (usually constructed by neural networks (NNs)) is essentially the generator $G_{\theta}(\cdot)$ in (2) [30]. During training, function $f^{-1}$ is learned to transform input data $\mathbf{x}$ (with condition $\mathbf{c}$) into $\mathbf{z}$ which follows a standard Gaussian distribution. Since the function $f^{-1}$ is invertible, once $f^{-1}$ is trained, its inverse $f$ is used to take random samples $\mathbf{z}$ (with condition $\mathbf{c}$) and generate $\mathbf{x}$. In flow-based models, function $f$ (or $f^{-1}$) is usually constructed by stacking multiple invertible transformations $f_{i}$, meaning $f=f_{1}\circ f_{2}...\circ f_{K}$ and $f^{-1}=f^{-1}_{K}\circ f^{-1}_{K-1}...\circ f^{-1}_{1}$. Fig 2 demonstrates how this stacked function transforms a simple Gaussian distribution into data distribution $p_{K}(\mathbf{x|c})$ and vice versa. By stacking invertible transformations $f_{i}$, the modeling capability of $f$ is also increased, enabling function $f$ to simulate more complex data distribution.
@@
Figure 1: Structure of flow-based models, where $\mathbf{x}$ and $\mathbf{\hat{x}}$ are the input data and generated data, $\mathbf{c}$ is the condition corresponding to data, $\mathbf{z}$ is a latent variable following Gaussian distribution, and $f$ is a bijective function $f:Z\rightarrow X$ which is usually constructed by NNs.
@@ III-B Combining Coupling Layer
$\mathbf{\hat{z}_{1}}=exp(s_{1}(\mathbf{x_{1};c}))\odot\mathbf{x_{2}}+t_{1}(\mathbf{x_{1};c})$ (9)
$\mathbf{(\hat{z}_{1},c)}=Combine(\mathbf{\hat{z}_{1},c})$ (10)
$\mathbf{\hat{z}_{2}}=exp(s_{2}(\mathbf{\hat{z}_{1};c}))\odot\mathbf{x_{1}}+t_{2}(\mathbf{\hat{z}_{1};c})$ (11)
$\mathbf{\hat{z}}=Combine(\mathbf{\hat{z}_{1},\hat{z}_{2}}),$ (12)
@@
Figure 3: The conditional transformation architecture of combining coupling layer $f_{ccl}$ and its inverse. $t$ and $s$ are the neural networks that can be expressed as $t(\cdot)$ and $s(\cdot)$, both $s$ and $t$ reduce the input dimensions as demonstrated in the figure. $\mathbf{x}$ and $\mathbf{\hat{x}}$ are the sampled and generated data, respectively. $\mathbf{z}$ and $\mathbf{\hat{z}}$ are the sampled and generated latent variables, respectively.
where $\mathbf{x}$ corresponds to the RLP data, the operation $Split(\cdot)$ partitions the input vector $\mathbf{x}$ (or $\mathbf{z}$) into two sub-vectors. $\mathbf{x_{1}}$ and $\mathbf{x_{2}}$, corresponding to the even and odd elements of $\mathbf{x}$, $s$ and $t$ are NNs, $Combine(\cdot)$ refers to the method used to merge two vectors into a single vector, the symbol $\odot$ indicates element-wise multiplication. The generation process of combining coupling layers $f_{ccl}$ can be expressed as
@@
$\mathbf{\hat{x}_{1}}=(\mathbf{z_{2}}-t_{2}(\mathbf{z_{1};c}))/exp(s_{2}(\mathbf{z_{1};c}))$ (15)
$\mathbf{(\hat{x}_{1},c)}=Combine(\mathbf{\hat{x}_{1},c})$ (16)
$\mathbf{\hat{x}_{2}}=(\mathbf{z_{1}}-t_{1}(\mathbf{\hat{x}_{1};c}))/exp(s_{1}(\mathbf{\hat{x}_{1};c}))$ (17)
$\mathbf{\hat{x}}=Combine(\mathbf{\hat{x}_{1},\hat{x}_{2}}).$ (18)
@@ IV Proposed Model: Full Convolutional Profile Flow
Figure 4: FCPFlow architecture is structured to process data through a series of FCPFlow blocks. $\mathbf{x}$ and $\mathbf{\hat{x}}$ are the sampled and generated RLPs, respectively. $\mathbf{z_{0}}$ and $\mathbf{\hat{z}_{0}}$ are the sampled and generated latent variables, respectively. $\beta$ and $\gamma$ are parameters in $f_{norm_{i}}$ introduced in Sec. IV-A, $W$ is a matrix that linearly transforms the input vector into another vector, $W$ is introduced in Sec. IV-C.
@@ IV-C Maximum Likelihood Estimation of FCPFlow
NL 60 minute 27,757
AUS[35] 30 minute 10,000
UK 30 minute 10,000
USA[36] 15 minute 9,110
UK weather[37] 30 minute 10,000
@@
NL 60 minute 365
UK 30 minute 365
USA[36] 15 minute 365
@@ VII Simulation Results for RLP Probabilistic Prediction
NL dataset 60 minute resolution
@@
UK dataset 30 minute resolution
@@
USA dataset 15 minute resolution